In how my ways can the letters of TAGORE be groued keeping in mind that the GOR are always together?

Grade:9

2 Answers

Utkarsh Gupta
14 years ago

TAGORE -- Let us take the letters GOR as one unit.

We have five blanks to fill --  _ _ _ _ _

No. of Permutations for the five spaces = 5! = 120.

No. of permutations for the GOR = 3! = 6

Total no. of permutations = 120*6 = 720(Ans.)

Gireesh Chandra Joshi
14 years ago

 

TAGORE -- Let us take the letters GOR as one unit.

We have five blanks to fill --  _ _ _ _

No. of Permutations for the five spaces = 4! = 24.

No. of permutations for the GOR = 3! = 6

Total no. of permutations = 24*6 = 144(Ans.)